1/2

William Wordsworth,  William Wordsworth,
Dafodiloj  The Daffodils

übersetzt von Krys Ungar

Mi vagis sola kiel nub’ Altflosas super la kampar’ Kaj tiam tuj aperis trup’ De dafodiloj; ora ar’ Apud la lag’, sub arba tend’ Flirtis kaj dancis en la vent’.

I wander’d lonely as a cloud

That floats on high o’er vales and hills, When all at once I saw a crowd, A host of golden daffodils, Beside the lake, beneath the trees Fluttering and dancing in the breeze.

Senbreĉe kiel astra bril’ Sur Lakta Voj’ flagretas, jen Sterniĝis ili en defil’ Senfine laŭ la laghaven’. Rigardon kaptis la balanc’ De dekmil kapoj en la danc’.

Continuous as the stars that shine And twinkle on the milky way, They stretch’d in never-ending line Along the margin of a bay: Ten thousand saw I at a glance Tossing their heads in sprightly dance.

Apudaj ondoj dancis, sed Triumfon gajnis floroj ĉi, Devigis gajon de poet’ Plej joviala kompani’: Mi rigardadis sen atent’, Ke min riĉigos la prezent’.

The waves beside them danced, but they Out-did the sparkling waves in glee: - A poet could not but be gay In such a jocund company: I gazed - and gazed - but little thought What wealth the show to me had brought.

Ĉar ofte, kiam kuŝas mi Penseme, aŭ en menslangvor’, Ekfulmas ili tra l’ konsci’ - La beno de solema hor’; La koro pro pleniga ĝu’ Kun dafodiloj dancas plu.

For oft, when on my couch I lie In vacant or in pensive mood, They flash upon that inward eye Which is the bliss of solitude; And then my heart with pleasure fills And dances with the daffodils.

…  …

2/2

Übersetzung des englischen Gedichtes “The Daffodils” von William Wordsworth (∗1770 – †1850) ins Esperanto durch Krys Ungar (Krys Bottrill).

Arg-39-632 (2009-09-20 15:01:36)

Verfasser dieses englischen Gedichtes ist William Wordsworth (∗1770 – †1850).

Arg-39-73 (2003-10-17 22:02:21)

Origina anglalingva teksto de WILLIAM WORDSWORTH (1770 - 1850)

http://www.poezio.net/version?poem-id=39&version-id=,632,73